Grade 12th passPhysical Chemistry

An aqueous solution containing MgCl2 and HCl was analysed by first titrating a 20.00 mL aliquot to a bromocresol green endpoint with 14.02 mL of 0.02932 M NaOH. A 10.00 mL aliquot was then diluted to 75.00 mL with distilled water and passed through a strong-acid ion exchange resin. The eluate and washings required 35.94 mL of the NaOH solution to reach the endpoint. Calculate the molar concentration of HCl and MgCl2 in the sample.

Jarra Moss
6 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
you have 20.00mL + 14.02mL = 34.02mL of solution of pH = 4.00 after titration
That will have
.. [H+] = 10^-4.00 = 1.000x10^-4 mol H+ / L
so mole [HCl] remaining after titration would be
.. 34.02mL * (1L / 1000mL) * (1.000x10^-4 mol HCl / L) = 3.402x10^-6 mol
moles HCl reacted would be 
.. 14.02 mL NaOH * (1L/1000mL) * (0.02932 mol/L) * (1mol HCl / 1 mol NaOH) = 4.1111x10^-4 
so moles HCl  present in 20.00mL aliquot is
.. 4.1111x10^-4 + 0.03402x10^-4 = 4.1451x10^-4
and that translates to a molarity of
.. (4.1451x10^-4 mol /  20.00mL) * (1000mL / L) = 0.020726 M HCl
********
let's discuss the strong acid resins on an side note... 
the strong acid resin contains SO3- groups which scavenge the Mg2+ and leave the two Cl- ions. 
The way the resins work is when they are fully regenerated, they are saturated with Na+ ions.  SO3- ---- Na+. 
 The Mg2+ ions displace the Na+ and are removed from the water stream.. SO3- ---- Mg2+ ----- SO3- Once the resins are saturated with Mg2+, they are regenerated by soaking them in brine which replaces the Mg2+ ions with Na+ and the process starts all over.
It is POSSIBLE to purchase new resins that are of the acid form and have H+ attached to the SO3- groups.  H+ instead of Na+.  We do it at my company.  But it's atypical and not cost effective.  Let's assume that's what's happening in this case so that
.. the Mg from MgCl2 gets adsorbed onto the resins and displaces H+
.. the H+ reacts with the Cl- and makes HCl which we test.
********** 
back to the problem
the 10.00 aliquot starts with
.. 10.00mL * (1L / 1000mL) * (0.020726 mol HCl / L) = 2.0726x10^-4 mol HCl
moles of HCl remaining after titration (same deal as above.. pH = 4.00) 
.. 75.00mL * (1L / 1000mL) * (1.000x10^-4 mol HCl / L) = 7.500x10^-6 mol
moles of HCl reacted with the NaOH during titration
.. 35.94mL NaOH * (1L/1000mL) * (0.02932 mol/L) * (1mol HCl / 1 mol NaOH) = 1.0538x10^-3
and we can write
.. mole HCl from acid + mol HCl from MgCl2 = mol HCl reacted + moles HCl remaining
and then 
.. mol HCl from MgCl2 = (1.0538x10^-3 + 7.500x10^-6) - 2.0726x10^-4 = 8.5404x10^-4
and mol of MgCl2 
.. 8.5404x10^-4 mol HCl * (1 mol MgCl2 / 2 mol HCl) = 4.2702x10^-4 mol MgCl2
and that is in a 10.00mL aliquot so that
.. (4.2702x10^-4 mol MgCl2 / 10.00mL) * (1000mL / 1L) = 4.2702x10^-2 M MgCl2
**********
notice I carry 1 extra sig fig during intermediate steps?  that's to avoid rounding errors
FINAL answer is
.. molarity HCl = 0.02073M
.. molarity MgCl2 = 0.04270M